OpenStudy (18dsears):

You are paid time-and-a-half for each hour worked over 40 hours a week. Last week you worked 50 hours and earned $660. What is your normal hourly salary? HELP!! I don't understand this..

OpenStudy (jdoe0001):

well last week you worked 50 hours that is, 40 regular plus 10 overtime for the first 40 hours, you got regular pay, let's say "x" is the rate for the last 10 hours, you got x + x/2, or time and a half so regular pay plus half that, or x + x/2 so, if we add those two together, that was $660 meaning \(\bf 40\cdot {\color{brown}{ x}}+10\cdot \left( {\color{brown}{ x}}+\cfrac{{\color{brown}{ x}}}{2} \right)=660\) solve for "x" :)

OpenStudy (18dsears):

thank you so much! this helps a lot;)